OpenStudy (anonymous):

C(7, 1) and D(4, -6) Distance =

OpenStudy (anonymous):

\[d(C, D) = \sqrt{(C_1 - D_1)^2 + (C_2 - D_2)^2} = \sqrt{3^2 + 7^2} = \sqrt{9 + 49} = \sqrt{58}\]

OpenStudy (aroub):

\[\sqrt{(-6-1)^{2}+(4-7)^{2}} = \sqrt{(-7)^{2}+(3)^{2}} =\sqrt{49 + 9} =\sqrt{58}\]